OpenStudy (anonymous):

y = c1e^x + c2e^−x is a family of solutions for the DE y′′ − y = 0. Find a solution with initial conditions: y(1) = 0, and y′(1) = e.

OpenStudy (anonymous):

how do i go about doing this

OpenStudy (anonymous):

i took intergrals 2 times and evaluated the family of equations

OpenStudy (anonymous):

but i could not get 0 or e

OpenStudy (anonymous):

y(1)=c1e^(1)+c2e^(-1)=0 ---> Eq. (1) y'(1)=c1e^1-c2e^(-1)=e ---> Eq. (2) solve for c1 & c2: Eq.(1)+Eq. (2) 2c1 e = e c1 = 1/2, plugin to either Eq (1) or Eq. (2), solve for c2 Eq. (1): (1/2)e+c2 e^(-1) =0 c2=-(1/2)e^2

OpenStudy (anonymous):

Do u hve any question on the solution here?

OpenStudy (anonymous):

i mean it gets the right answer but i think i am messing up with the integrarion of ce^x

OpenStudy (anonymous):

hmm you dont integrate here just solve for c1 and c2 first

OpenStudy (anonymous):

i guess its fine then

OpenStudy (anonymous):

y = c1e^x + c2e^−x y =(1/2)e^x + (-e^2/2)e^−x solution to DE

OpenStudy (anonymous):

ok good luck now, have fun solving ..lol